# DualNILM: Energy Injection Identification Enabled Disaggregation with Deep Multi-Task Learning

### v1

@@ IV-A Method Overview and Design Rationale
Multi-Scale Feature Extraction: Appliance usage patterns often manifest as short-term events (e.g., on/off transients) superimposed on longer-term trends, whereas injection can exhibit relatively smooth variations (e.g., daily solar generation). To capture both fine-grained patterns and global context, DualNILM combines convolutional layers (for local feature learning) with Transformer layers (for longer-range sequence dependencies).
@@ IV-B Model Architecture
The overall architecture of DualNILM is illustrated in Fig. 2. Our model comprises four main components: 1) CNN encoders, 2) transformer encoders, 3) a transformer decoder, and 4) task-specific projection layers. This design allows for the effective capture of both local and global temporal dependencies in the power consumption patterns while enabling multi-task learning.
IV-B1 CNN Encoders
@@ IV-B2 Transformer Encoder
IV-B3 Transformer Decoder and Task-Specific Projections
@@
IV-C Highlights of DualNILM
Robust in Real Settings: Apply LayerNorm [32] instead of normalization for inputs to ensure numerical stability when faced with unscaled, continuous data. In practice, DualNILM can process raw meter readings without mandatory preprocessing, mitigating performance drops due to distributional shifts between training and deployment.
Scalable & Incremental: Each appliance’s state recognition is handled by its own projection layer, allowing new appliances to be added without retraining the entire network. This modular “add-on” structure preserves performance for existing appliances and only requires fitting new parameters for the introduced appliance.
Cross-Task Synergy: Jointly modeling appliance states and energy injection helps disentangle negative power flows from reduced consumption. Accurate inverter detection can prevent misclassifying low-usage intervals as “off,” while robust appliance recognition clarifies whether dips in aggregate load stem from actual load decrease or solar export.
Overall, DualNILM integrates CNN-based local feature extraction, transformer-level global context, and dual-task outputs to tackle behind-the-meter injection challenges with stable performance, easy expandability, and improved separation of negative flows.
IV-D Training Objective and Optimization
@@
where and are weighting factors to balance the two tasks. The model can be optimized via gradient backpropagation approaches like SGD [33], Adam [34], etc. Specifically, we employ the Dice loss [35], which is particularly effective for imbalanced binary classification problems for appliance state recognition:
For injection estimation, we can use a standard L1 or L2 measure: MAE (Mean Absolute Error) and Mean Square Error (MSE). Note that we can apply constraints/normalization on injection disaggregation targets as the maximum output of micro-inverters or community solar installations often adheres to standardized designed values like EN50583 [36].
@@ V-A Datasets and Simulation
V-A1 Laboratory Dataset
We collected a real-world dataset in our laboratory and conducted extensive experiments. The dataset includes individual power consumption data for multiple appliances and the behind-the-meter energy injection from a micro-inverter. The data was collected from December 6 to December 15, 2023, and was sampled at 2-second intervals using sub-meters for each appliance on both active power and reactive power. The dataset includes the following appliances: i) Air purifier, ii) Heater, iii) Light bulb Type I, iv) Light bulb Type II, v) Air compressor, vi) Air conditioner, vii) Micro-inverter (energy injection). Table. I presents the key statistics of our laboratory dataset.
@@ V-A1 Laboratory Dataset
V-A2 Synthetic Public Datasets
To evaluate the generalizability of DualNILM and facilitate comparison with existing methods, we simulate energy injection scenarios using two widely-used public NILM datasets: REDD [13] and UK-DALE [37]. Our simulation leverages real-world solar irradiance data and photovoltaic (PV) system modeling to create realistic scenarios of residential solar energy injection. The simulation process involved the following steps:
@@ V-A2 Synthetic Public Datasets
Simulation of Photovoltaic System Output: Using the collected irradiance data, we simulate a residential PV system with a designed maximum output capacity of 120 W, representing a small-scale household solar installation. The simulation accounts for temperature effects on PV efficiency, following the methods described in [39, 40].
@@
Figure 3 demonstrates the challenging disaggregation scenario on UKDALE House 1 during midday hours (10:30-13:30) on April 05, 2013, when PV injection interacts with typical household consumption. The aggregate power exhibits dramatic variations as appliances cycle, yet the PV injection (green shaded area) creates subtle but critical distortions in the net measurement. During periods of low baseline consumption, such as when only the refrigerator operates between switching events, even modest PV generation can mask appliance signatures entirely.
@@ V-B Evaluation Metrics
We employ different metrics to evaluate the performance of DualNILM on appliance state recognition and energy injection estimation tasks.
V-B1 Appliance State Recognition Metrics
For each appliance and the overall model performance, we compute the following metrics:
Accuracy: The proportion of correct predictions (both true positives and true negatives) among the total number of cases examined.
Recall: The proportion of actual positive cases that were correctly identified.
Precision: The proportion of positive identifications that were actually correct.
F1-score: The harmonic mean of precision and recall, providing a single score that balances both metrics.
These metrics are calculated for each appliance individually and then averaged across all appliances to provide an overall performance assessment. Noted that since the
V-B2 Energy Injection Estimation Metric
@@ V-C Benchmark Methods
For fair comparison and evaluation, we select both traditional NILM methods like FHMM and modern deep learning approaches. While some methods specialize in either appliance state recognition or energy injection disaggregation, methods like FHMM and UNetNILM handle both tasks simultaneously. This diverse selection enables comprehensive evaluation of DualNILM’s performance across different technical approaches and computational paradigms, with particular attention to the unique challenges posed by behind-the-meter energy injection.
Detailed configurations and implementation specifics are provided in Appendix D-A. All benchmark models were implemented with carefully tuned hyperparameters to ensure fair comparison. This comprehensive selection enables evaluation across different technical approaches, with particular attention to methods capable of handling behind-the-meter energy injection.
@@ VI Experimental Results
This section presents comprehensive evaluation of DualNILM across real-world laboratory measurements and synthetic public datasets augmented with PV injection. We analyze performance on controlled micro-inverter injection collected in our laboratory and validate generalization through synthetic augmentation of REDD and UK-DALE datasets, demonstrating the framework’s effectiveness across diverse behind-the-meter injection scenarios.
@@
| Laboratory Dataset (Real Behind-the-Meter Injection) | FHMM | 60.95 | 28.32 | 16.84 | 17.25 | |||
@@
VI-A Overall Performance Overview
Tables III and IV report comprehensive performance metrics for appliance-state recognition and energy-injection disaggregation. We evaluate under two complementary settings: i) controlled laboratory experiments using a real 500 W micro-inverter as the behind-the-meter injection; and ii) household datasets (REDD and UK-DALE) augmented with synthetic PV profiles to emulate typical residential consumption–generation patterns. This multi-faceted design enables robust validation across diverse, realistic household usage regimes.
VI-B Performance on Laboratory Dataset
@@
Among deep learning approaches, we observe a clear performance gradient correlating with architectural complexity. Seq2Point, designed for point-wise load disaggregation, achieves 34.27% F1-score as its narrow temporal window cannot distinguish between appliance transitions and injection variations. CNN-LSTM improves to 36.67% by incorporating temporal memory, while the Transformer reaches 42.76% through self-attention mechanisms that capture longer-range dependencies. UNetNILM, with its U-Net architecture originally designed for multi-task NILM, achieves 53.89% F1-score by jointly modeling states and power consumption, though it lacks explicit injection modeling.
DualNILM achieves 84.54% F1-score, representing a 57% relative improvement over the best baseline. This dramatic improvement validates our dual-task hypothesis: explicitly modeling injection as a parallel objective enables the network to disambiguate genuine appliance activities from generation-induced variations. Figure 4 visualizes this capability, showing accurate state detection even during concurrent inverter operation periods.
@@
For continuous injection estimation (Table IV), the laboratory results reveal interesting trade-offs between different regression approaches. Seq2Seq achieves competitive MAE (0.2622) through its sequence-to-sequence architecture optimized for time series prediction. However, its higher RMSE (0.3512) indicates susceptibility to outliers during rapid injection transitions. DAE shows the poorest performance (RMSE: 0.8830), as its denoising objective is misaligned with the injection estimation task. UNetNILM demonstrates reasonable performance (RMSE: 0.2305) by leveraging its multi-scale feature extraction.
DualNILM achieves the best overall performance with RMSE of 0.1429 and MAE of 0.0804. The superior accuracy stems from the synergistic relationship between tasks: accurate appliance state detection helps refine injection estimates by accounting for consumption changes, while precise injection estimation prevents false appliance detections during generation fluctuations.
| Laboratory Dataset (Real Behind-the-Meter Injection) | FHMM | 0.5015 | 0.4492 | |||||
@@
| REDD (Synthetic PV Injection) | House 1 | FHMM | 0.8093 | 0.2200 | ||||
@@
| UKDALE (Synthetic PV Injection) | House 1 | FHMM | 1.1905 | 0.3331 | ||||
@@
VI-C Performance on Synthetic Public Datasets
@@
Traditional NILM Approaches.
@@
XGBoost, as a gradient boosting ensemble method, shows interesting dataset-dependent behavior. On REDD, it maintains reasonable performance (F1: 79.18% House 1, 93.53% House 2) through handcrafted features including time-of-day, power derivatives, and statistical aggregates. However, it degrades significantly on UK-DALE (F1: 45.97% House 1, 70.67% House 2), suggesting that its feature engineering is less effective for UK appliance signatures and usage patterns. This highlights a fundamental limitation of traditional machine learning: the need for manual feature design that may not generalize across different geographical contexts.
@@
Deep learning methods exhibit varied resilience to injection based on their architectural inductive biases:
Seq2Point maintains high accuracy () but suffers from precision-recall imbalance. Its point-wise prediction cannot leverage temporal context to distinguish injection from load reduction, resulting in conservative predictions biased toward ”off” states (F1: 36-50%).
@@
Transformer demonstrates improved resilience (F1: 30-61%) through self-attention, which can correlate time-of-day with typical solar generation patterns. However, without explicit injection modeling, it still confuses cloud-induced generation drops with appliance activations.
UNetNILM exhibits high variance across houses, from 57.10% (REDD House 3) to 75.27% (REDD House 1). This instability suggests difficulty balancing its dual objectives of state classification and power regression when injection is present but not explicitly modeled.
@@
DualNILM maintains robust performance across all test scenarios. On REDD, it achieves F1-scores of 96.64%, 93.11%, and 80.04% for Houses 1-3 respectively. On UK-DALE, it achieves 94.71% and 92.31% for Houses 1-2. This consistency across geographically diverse datasets with different appliance types, usage patterns, and synthetic injection profiles demonstrates genuine generalization capability.
The slightly lower performance on REDD House 3 (80.04%) provides valuable insights. This house contains more low-power appliances whose signatures are easily masked by injection variations. Yet DualNILM still substantially outperforms all baselines, validating the robustness of the dual-task approach even in challenging scenarios.
@@
Table IV reveals that injection estimation becomes more challenging on public datasets compared to laboratory settings. This is expected as the synthetic injection exhibits more complex temporal dynamics including cloud transients and seasonal variations. Seq2Seq occasionally achieves competitive RMSE on individual houses but shows less consistent MAE performance. Notably, UNetNILM exhibits severe instability on UK-DALE House 1 (RMSE: 7.5014), suggesting training convergence issues when the injection magnitude varies widely.
@@ VI-D Key Insights and Analysis
The performance gap between DualNILM and UNetNILM, both multi-task architectures, underscores the importance of explicitly modeling injection rather than treating it as a general regression target. While UNetNILM jointly performs classification and regression, it lacks the architectural separation that enables DualNILM to learn distinct representations for consumption and generation.
@@
The laboratory experiments capture real hardware effects including inverter switching noise, electromagnetic interference, and actual appliance transients that are difficult to simulate. Conversely, the synthetic datasets provide diversity in household types, appliance portfolios, and usage patterns essential for evaluating generalization. Together, they provide comprehensive validation that neither alone could achieve.
Method-Specific Failure Modes.
Traditional methods (FHMM) fail due to violated mathematical assumptions. Machine learning methods (XGBoost) struggle with feature engineering that doesn’t generalize. Deep learning methods without temporal modeling (Seq2Point) cannot disambiguate temporal patterns. Even sophisticated architectures (Transformer, UNetNILM) fail without explicit injection modeling. These diverse failure modes validate that the injection challenge requires fundamental architectural innovation, not incremental improvements.
@@
The results have immediate practical implications for smart grid deployment. Existing NILM systems will experience severe degradation in solar-equipped homes, particularly during peak generation hours. The consistent performance of DualNILM across diverse scenarios suggests it could be deployed as a software upgrade to existing smart meter infrastructure, enabling continued load monitoring despite distributed generation.

### v2

@@ IV-A Design Rationale
Multiscale Feature Extraction: Appliance usage patterns often manifest as short-term events (e.g., ON/OFF transients) superimposed on longer-term trends, whereas injection can exhibit relatively smooth variations (e.g., daily solar generation). To capture both fine-grained patterns and global context, DualNILM combines convolutional layers (for local feature learning) with Transformer layers (for longer-range sequence dependencies).
@@ IV-B Model Architecture
The overall architecture of DualNILM is illustrated in Figure 2. Our model comprises four main components: 1) CNN encoders, 2) transformer encoders, 3) a transformer decoder, and 4) task-specific projection layers. This design allows for the effective capture of both local and global temporal dependencies in the power consumption patterns while enabling multi-task learning.
CNN Encoders.
@@
Transformer Decoder & Task-Specific Projections.
@@
For enhanced cross-task synergy, the final injection estimate incorporates the predicted inverter state through element-wise filtering:
(10) where denotes the Hadamard product. This cross-task constraint ensures injection predictions are zeroed when the inverter is predicted inactive, fundamentally improving the physical consistency of the dual-task outputs.
Design Justification for DualNILM
The proposed architecture addresses the signal eclipse phenomenon directly through the above coupled design choices, ensuring both theoretical soundness and practical viability. First, the CNN-Transformer hybrid accommodates distinct temporal characteristics: appliance switching events exhibit sharp transients captured by convolutional kernels, while energy injection exhibits complex multiscale temporal dependencies and cross-correlation patterns requiring long-range attention mechanisms to capture their dynamics. Second, explicit separation of injection estimation through dedicated Seq2Seq projection prevents mathematical contamination of appliance state classification, where negative flows would corrupt binary decision boundaries. This architectural separation, combined with LayerNorm stabilization [28], enables direct processing of raw meter readings while maintaining modular scalability for incremental appliance addition. Third, the shared encoder representations enable cross-task synergy that significantly reduces the problem of identification, allowing accurate disambiguation between actual load reduction and injection-induced power dips within their respective constraint domains.
IV-C Training Objective and Optimization
@@
where is the weighting factor to balance the dual targets.
Our model can be optimized using gradient backpropagation approaches such as SGD [29], Adam [30], etc. Specifically, we employ Dice loss [31], which is particularly effective for imbalanced binary classification problems for appliance state recognition:
For injection estimation loss, we use the standard L2 measure: Root Mean Square Error (RMSE). Note that we can apply constraints/normalization on injection disaggregation targets as the maximum output of micro-inverters or community solar installations typically must adhere to standardized values like EN50583 [32].
@@ V-A Datasets and Simulation
Laboratory Dataset
We collected a real-world dataset in our laboratory and conducted extensive experiments. The dataset includes individual power consumption data for multiple appliances and the BTM energy injection from a micro-inverter. The data was collected from December 6 to December 15, 2023, and was sampled at 2-second intervals using sub-meters for each appliance on both active power and reactive power. The dataset includes the following appliances: i) Air purifier, ii) Heater, iii) Light bulb Type I, iv) Light bulb Type II, v) Air compressor, vi) Air conditioner, vii) Micro-inverter (energy injection). Table I presents the key statistics of our laboratory dataset.
@@
Synthetic Public Datasets
To evaluate the generalizability of DualNILM and facilitate comparison with existing methods, we simulate energy injection scenarios using two widely-used public NILM datasets: REDD [9] and UK-DALE [33]. Our simulation takes advantage of real-world solar irradiance data and photovoltaic (PV) system modeling to create realistic scenarios of residential solar energy injection. Detailed implementation will be open-source and available after acceptance. The simulation process involved the following steps:
@@
Simulation of Photovoltaic System Output: Using the collected irradiance data, we simulate a residential PV system with a designed maximum output capacity of 120 W, representing a small-scale household solar installation. The simulation accounts for the effects of temperature on the efficiency of PV, following the methods described in [35, 36].
@@
Figure 3 shows an example to demonstrate a challenging disaggregation scenario on UKDALE House 1 during the midday hours (10:30-13:30) on April 05, 2013, when PV injection interacts with typical household consumption. The aggregate power exhibits dramatic variations as appliances cycle, yet the PV injection (green shaded area) creates subtle but critical distortions in the net measurement. During periods of low baseline consumption, such as when only the refrigerator operates between switching events, even modest PV generation can mask appliance signatures entirely.
@@ V-B Evaluation Metrics
Appliance State Recognition Metrics
For appliance state recognition tasks, we report Accuracy, Recall, Precision, and F1-score. Since the imbalanced nature of the appliance ON/OFF state, the F1-score serves as the most comprehensive metric.
All above metrics are calculated for each appliance individually and then averaged across all appliances to provide an overall performance assessment.
Energy Injection Estimation Metric
@@ V-C Benchmark Methods
For fair comparison and evaluation, we select both traditional NILM methods, such as FHMM, and modern deep learning approaches. Although some methods specialize in appliance state recognition or energy injection disaggregation, methods such as FHMM and UNetNILM handle both tasks simultaneously. This diverse selection enables comprehensive evaluation of DualNILM’s performance across different technical approaches and computational paradigms, with particular attention to the unique challenges posed by BTM energy injection.
Detailed configurations and implementation specifics are provided in Appendix D-A. All benchmark models were implemented with carefully tuned hyperparameters to ensure fair comparison. This comprehensive selection enables evaluation of different technical approaches, with particular attention to methods capable of handling BTM energy injection.
@@ VI Experimental Results
This section presents a comprehensive evaluation of DualNILM across real-world laboratory measurements and synthetic public datasets augmented with PV injection. We analyze performance on controlled micro-inverter injection collected in our laboratory and validate generalization through synthetic augmentation of REDD and UK-DALE datasets, demonstrating the framework’s effectiveness across diverse BTM injection scenarios.
@@
| Laboratory Dataset (Real Injection) | FHMM | 60.95 | 28.32 | 16.84 | 17.25 | |||
@@
| REDD with Synthetic Injection | House 1 | FHMM | 79.52 | 1.46 | 13.47 | 2.07 | ||
@@
| UKDALE with Synthetic Injection | House 1 | FHMM | 89.25 | 1.58 | 0.76 | 1.01 | ||
@@
Overview.
Tables III and IV report comprehensive performance metrics for appliance-state recognition and energy-injection disaggregation. We evaluated them in two complementary settings: i) controlled laboratory experiments using a real 500 W micro-inverter as the BTM injection; and ii) household datasets (REDD and UK-DALE) augmented with synthetic PV profiles to emulate typical residential consumption–generation patterns. This multi-faceted design enables robust validation across diverse, realistic household usage regimes.
VI-A Performance on Laboratory Dataset
@@
Among deep learning approaches, we observe a clear performance gradient correlating with architectural complexity. Seq2Point, designed for point-wise load disaggregation, achieves an average score of 34.27% F1 as its narrow temporal window cannot distinguish between appliance transitions and injection variations. CNN-LSTM improves to 36.67% by incorporating temporal memory, while the Transformer reaches 42.76% through self-attention mechanisms that capture longer-range dependencies. UNetNILM, with its U-Net architecture originally designed for multi-task NILM, achieves 53.89% F1-score by jointly modeling states and power consumption, though it lacks explicit injection modeling.
DualNILM achieves 84.54% F1-score, representing a 57% relative improvement over the best baseline. This dramatic improvement validates our dual-task hypothesis: explicitly modeling injection as a parallel objective enables the network to disambiguate genuine appliance activities from generation-induced variations. Notably, XGBoost achieves 47.75% F1-score through its tree-based ensemble learning, which implicitly discovers patterns in the dual-channel input. However, unlike DualNILM, it cannot provide explicit injection estimates, and its performance varies significantly across different scenarios. Figure 4 visualizes this capability, showing accurate state detection even during concurrent inverter operation periods.
@@
For continuous injection estimation (Table IV), the laboratory results reveal interesting trade-offs between different regression approaches. Seq2Seq achieves competitive MAE (0.2622) through its sequence-to-sequence architecture optimized for time series prediction. However, its higher RMSE (0.3512) indicates susceptibility to outliers during rapid injection transitions. DAE shows the poorest performance (RMSE: 0.8830), as its denoising objective is misaligned with the injection estimation task. UNetNILM demonstrates reasonable performance (RMSE: 0.2305) by leveraging its multiscale feature extraction.
DualNILM achieves superior overall performance, yielding average RMSE and MAE of 0.1429 and 0.0804, respectively. The superior accuracy stems from the synergistic relationship between tasks: accurate appliance state detection helps refine injection estimates by accounting for consumption changes, while precise injection estimation prevents false appliance detections during generation fluctuations.
| Laboratory Dataset (Real Injection) | FHMM | 0.5015 | 0.4492 | |||||
@@
| REDD with Synthetic Injection | House 1 | FHMM | 0.8093 | 0.2200 | ||||
@@
| UKDALE with Synthetic Injection | House 1 | FHMM | 1.1905 | 0.3331 | ||||
@@
VI-B Performance on Synthetic Public Datasets
@@
Performance of Traditional NILM Approaches.
@@
XGBoost, as a gradient boosting ensemble method, shows interesting dataset-dependent behavior. On REDD, it maintains reasonable performance (F1-scores: 79.18% House 1, 93.53% House 2) through handcrafted features including time-of-day, power derivatives, and statistical aggregates. However, it degrades significantly on UK-DALE (F1-scores: 45.97% House 1, 70.67% House 2), indicating sensitivity to dataset-specific appliance distributions and consumption patterns. This highlights a fundamental limitation of traditional machine learning: the need for manual feature design that may not generalize across different geographical contexts.
@@
Deep learning methods exhibit varied resilience to PV injection based on their architectural inductive biases:
Seq2Point maintains high accuracy () but suffers from precision-recall imbalance. Its point-wise prediction cannot leverage temporal context to distinguish injection from load reduction, resulting in conservative predictions biased toward OFF states (F1-scores: 36-50%).
@@
Transformer demonstrates improved resilience (F1-scores: 30-61%) through self-attention, which can correlate time-of-day with typical solar generation patterns. However, without explicit injection modeling, it still confuses cloud-induced generation drops with appliance activations.
UNetNILM exhibits high variance across houses, F1-scores from 57.10% (REDD House 3) to 75.27% (REDD House 1). This instability suggests difficulty balancing its dual objectives of state classification and power regression when injection is present but not explicitly modeled.
@@
DualNILM maintains robust performance across all test scenarios. On REDD, it achieves F1-scores of 96.64%, 93.11%, and 80.04% for Houses 1-3, respectively. On UK-DALE, it achieves 94.71% and 92.31% for Houses 1-2. This consistency across geographically diverse datasets with different appliance types, usage patterns, and synthetic injection profiles demonstrates genuine generalization capability.
The slightly lower performance on REDD House 3 (F1-score 80.04%) provides valuable insights. This house contains more low-power appliances whose signatures are easily masked by injection variations. Yet DualNILM still substantially outperforms all baselines, validating the robustness of the dual-task approach even in challenging scenarios. The performance variance of XGBoost across REDD houses (50.51% to 93.53% F1-score) and its degraded performance on UK-DALE (45.97% on House 1) suggest that tree-based pattern learning, while effective in specific scenarios, lacks the generalization capability of explicit multi-task modeling.
@@
Table IV reveals that injection estimation becomes more challenging on public datasets compared to laboratory settings. This is expected as the synthetic injection exhibits more complex temporal dynamics including cloud transients and seasonal variations. Seq2Seq occasionally achieves competitive RMSE on individual houses but shows less consistent MAE performance. Notably, UNetNILM exhibits severe instability on UK-DALE House 1 (RMSE: 7.5014), suggesting training convergence issues when the injection magnitude varies widely. This instability, contrasted with its reasonable performance on REDD datasets, highlights the sensitivity of standard multi-task architectures to dataset-specific characteristics without proper architectural design for handling bidirectional power flows.
@@ VI-C Key Insights and Analysis
The performance gap between DualNILM and UNetNILM, both multi-task architectures, underscores the importance of explicitly modeling injection rather than treating it as a general regression target. Although UNetNILM jointly performs classification and regression, it lacks the architectural separation and cross-task synergy that enables DualNILM to learn different, implicitly connected representations for consumption and generation.
@@
Laboratory experiments capture real hardware effects, including inverter switching noise, electromagnetic interference, and actual appliance transients that are difficult to simulate. In contrast, synthetic datasets provide diversity in household types, appliance portfolios, and usage patterns essential for evaluating generalization. Together, they provide comprehensive validation that neither could alone achieve.
Method-Specific Failure Modes and Optimization Challenges.
Our results reveal distinct failure patterns that illuminate the fundamental challenges of BTM injection.
Traditional methods, such as FHMM, fail due to violated mathematical assumptions, and their factorial structure assumes discrete states incompatible with continuous injection variations. Notably, XGBoost serves as a strong baseline that achieves competitive performance (74.41% average F1 on REDD from Table XIII) through implicit feature engineering via recursive tree splits on the dual-channel input. However, it is fundamentally limited by its inability to provide explicit injection estimates and its reliance on pattern memorization, which makes it susceptible to distributional shifts.
Deep learning methods encounter more complex optimization challenges. The introduction of the reactive power channel and multiple appliance types with imbalanced activation patterns significantly expands the optimization landscape. Seq2Point (34.27% F1 on laboratory data) cannot disambiguate temporal patterns due to its narrow window. CNN-LSTM and Transformer architectures, despite sophisticated temporal modeling, struggle with convergence since the simultaneous presence of sparse appliance activations and continuous injection creates conflicting gradient signals. These diverse failure modes validate that the injection challenge requires fundamental architectural innovation, not incremental improvements.
@@
The results have immediate practical implications for the deployment of smart grids. Existing NILM systems will experience severe degradation in solar or BTM generation-equipped homes, particularly during peak generation hours. Our analysis shows that performance degradation is not uniform but depends critically on the relationship between injection magnitude and base load levels. When injection exceeds 40% of typical consumption, even sophisticated baselines fail to maintain acceptable accuracy. The consistent performance of DualNILM across diverse scenarios suggests it could be deployed as a software upgrade to existing smart meter infrastructure, enabling continued load monitoring despite distributed generation.